7. Multiple regression analysis is used to study how an individuals income (y, in thousands of dollars) is influenced by age(c) Perform a t test to determine whether the coefficient for the level of education is significantly different from zero.
Answer #1

a)

SSexplained = 84

SS residual = 112

SStotal = SSexplained + SS residual =84 + 112 = 196

Multiple co-efficient of determination = 1- SS residual / SStotal = 1- 112 / 196 = .4286

It means with the help of 3 regressor variable we are able to explain 42.86 % of variability in individual's income.

b)

If co-efficient of x3 is significant then it means when x1 and remain fixed then x3= 1 i.e a male tends to income 510 ( .51 *1000) dollars less than when x3=0 i.e female. In other words the income of a female is 510 more than a male with same age and level of education.

But if we do t test then it is seen t = -.52/ .92 = - 0.55

tcrit = t \small \alpha /2 =0.025,18 = 2.101 > | t |

So reject the null hypothesis. Gender does not have effect on income. So males and females tend to income same when age & level of education are same.

c)

H0:  \small \beta2 = 0

H1:  \small \beta2 \small \neq 0

\small \hat{\beta }= .92

SE(\small \hat{\beta }) =.19

t = \small \hat{\beta } / SE(\small \hat{\beta }) = 4.84

tcrit = t \small \alpha /2 =0.025,18 = 2.101 < | t |

So, we fail to reject null hypothesis.

So, level of education is significantly different from 0.

Feel free to comment if you need any clarification.

If you find my answer useful please put thumbs up. Thank you.